OpenStudy (anonymous):

find sn, fo the given arithmetic series a1-10, d=-9, n=15

OpenStudy (anonymous):

is that a1 = 10 or -10?

OpenStudy (anonymous):

Well, anyway... I'm going to assume its -10 The sum of an arithmetic series: n Sn = -- (A1 + An) 2 An = A1 + d (n-1) = -10 + -9(n-1) = -10 -9n +9 = -9n - 1 So A15 = -9(15) -1 = -136 15 Sn = --- (-10 + -136) = -1095 2

OpenStudy (anonymous):

@christiner14